OpenStudy (anonymous):

p - 5 - 13p = 150 - 11

OpenStudy (swissgirl):

First we gotta clean up and gather like terms

OpenStudy (swissgirl):

we gotta do a lil of adding and subtracting on either side.

OpenStudy (swissgirl):

Do you have any idea?

OpenStudy (anonymous):

here is the original equation: \[p-5-13p=150-11\]first step is to simplify the math on the right simply by taking 150 and subtracting 11. Then you get:\[p-5-13p=139\]next step is to add 5 to both sides of the equation giving you:\[p-13p=139+5\]and that simplifies to:\[p-13p=144\] next you add the left side of the problem together. It might be easier to think of it like this:\[1p-13p=144\]which simplifies to:\[-12p=144\] the last step is to divide by -12 on both sides:\[p=144/-12\]which means that p is equal to -12. This can easily be checked by substituting -12 for p in the original equation.\[-12-5-13(-12)=150-11\]\[-12-5+156=139\]\[-17+156=139\]\[139=139\]

OpenStudy (anonymous):

150 - 10 = 140 sp find out what numbers you can put in the variables (the letters) to make it equal to 140... I HOPE im right.. im not a teacher.